### v1

@@ Abstract
Similarly, the sliding mode frequency exhibits a clear anti-correlation with the sliding amplitude (compare red and brown traces in Fig. 4a): as the amplitude increases, the frequency softens. Since both the lattice temperature and the level of electronic excitation are fixed in double-pulse experiments, a significant contribution from electronic softening, as reported in other semimetals [52], can be excluded. We therefore attribute the observed amplitude-dependent frequency shift to intrinsic lattice anharmonicity, explored by large-amplitude coherent nuclear motion.
We note that these insights into the processes underlying coherent phonon decay cannot be straightforwardly obtained from conventional (fluence-dependent) optical pump–probe experiments alone, but require precise control over the phonon amplitude in a targeted excited state, as enabled by double-pulse excitation (compare pump-probe and pump-pump-probe data in Fig. 4b)).
In a last step, we consider the impact of large-amplitude sliding motion on the electronic structure. For the highest sliding amplitudes in the excited state, Fourier transforms of the time-resolved SHG traces show an additional contribution at twice the sliding mode frequency ($2\omega_{\text{s}}$, see Fig. 5). This component is absent for single-pulse excitation, for double-pulse excitation near temporal overlap (Fig. 5a-c), and in transient reflectivity studies [33, 32, 35, 34]. This suggests that its generation requires vibrational amplitudes beyond the sliding mode saturation threshold, achievable only through sequential excitation. Phonon harmonics have been attributed to anharmonicities in the interatomic potential [53], or to periodic modulation of electronic states through electron-phonon coupling [37, 54, 55]. In the first case, frequency doubling originates directly from anharmonicities in the atomic motion $Q(t)$, which lead to a second harmonic component in the lattice displacement. In the second case, the coherent lattice motion periodically modulates the material’s second-order susceptibility, $\chi^{(2)}=\chi^{(2)}(t)$, thereby generating sidebands in the SHG spectrum through dynamic electron–phonon coupling. This mechanism is commonly associated with deformation potential effects [37] or phonon Floquet phenomena [56], where the electronic structure is periodically driven by vibrational motion.
Although second harmonic phonon generation in the intrinsically anharmonic potential of WTe2 may contribute, several experimental observations point to a modulation of $\chi^{(2)}$ as the most likely origin of the $2\omega_{\text{s}}$ feature. First, recent transient reflectivity measurements on the related compound Td-MoTe2 have reported similar harmonic and sum-frequency components for probe photon energies near electronic resonances [37]. Second, trARPES has uncovered periodic shifts of electronic band energies in WTe2 following sliding mode excitation [26], corroborated by DFT calculations [35]. Together, these results suggest a scenario in which amplified lattice motion periodically modulates the energy of electronic states, thereby driving the nonlinear probing process in and out of resonance. The resulting amplitude modulation of the optical SHG response leads to the emergence of frequency sidebands (see derivation in the Methods section).
While phonon-induced energy shifts near the Fermi level are well established, our results show that repeated optical excitation can modulate optical transitions even in band structure regions that are typically only weakly coupled to structural modes. This capability is particularly relevant for the generation of optical frequency combs via coherent phonons [54, 55]. Our work extends this approach to a broader range of photon energies, beyond linear transmission or reflectivity, and to the regime of optical second-harmonic generation, offering an alternative pathway to the modulation of optical nonlinearity that is complementary to recent advances in Floquet engineering [18].
@@
In summary, we have shown that redistributing the energy of a single ultrafast excitation into two or more pulses can overcome the amplitude limitation of key vibrational modes. Extending this concept to multi-pulse sequences of sub-threshold excitations may enable ’cold’ driving of Raman-active modes deep into the nonlinear regime [1, 15, 16]. By analogy with nonlinear phononics [3, 8, 9], the resulting large-amplitude motion along Raman-active coordinates could access novel dynamical phases of matter and facilitate ultrafast, energy-efficient switching of order parameters. Our approach may be particularly relevant for the class of sliding ferroelectrics, where—due to broken inversion symmetry—decisive modes are typically both Raman and IR-active but carry only small dipole moments, limiting the efficiency of direct resonant excitation.
While a linear relationship between electronic temperature and vibrational amplitude is a central assumption of the DECP model [12], there is no fundamental reason for it to hold universally. The principles demonstrated here for WTe2 should therefore be broadly applicable to other materials and their structural modes. In this context, high-amplitude vibrational spectroscopy enabled by sequential excitation [57, 46] offers a powerful tool to probe phonon–electron and phonon–phonon interactions in nonequilibrium states beyond conventional pump–probe schemes. Understanding how decisive coherent phonons decay under such conditions is essential for advancing vibrational control strategies. For WTe2, our work identifies a nonequilibrium decay channel directly governed by coherent nuclear motion, highlighting amplitude-dependent damping as a key factor in determining phonon lifetimes as well as energy and heat transfer. We envision that the enhancement of vibrational coherences near topologically trivial and paraelectric structural configurations opens new avenues for coherent control of both topology and ferroelectricity in functional materials.
@@ Methods
Sample preparation. WTe2 single crystals (commercially ordered from HQ Graphene) were exfoliated to remove the oxidized surface layers and immediately mounted in an optical microscopy cryostat. The sample space was constantly purged with nitrogen gas to minimize oxidization of the pristine surface layers. Peak SHG and oscillatory signals from the sample showed less than $2\,\%$ degradation over several weeks. All experiments were carried out at room temperature.
@@
Experimental setup. In single-pulse trSHG measurements, structural dynamics are initiated by focusing pump pulses (center wavelength $\lambda_{\text{p}}=1.5\,\mu\text{m}$, photon energy $\hbar\omega_{\text{p}}=0.83\,\text{eV}$, pulse duration $\Delta\tau_{\text{p}}=100\,\text{fs}$, absorbed fluence $F_{\text{p}}=0.4–6.5\,\text{mJ}\,\text{cm}^{-2}$, polarization along the $b$-axis of the WTe2 crystal, repetition rate $f_{\text{rep}}=1\,\text{kHz}$) onto the sample surface under normal incidence using a plano-convex lens ($f_{\text{pump}}=100\,\text{mm}$). The resulting dynamics are probed by monitoring the second harmonic signal generated by probe pulses (center wavelength $\lambda_{\text{pr}}=0.8\,\mu\text{m}$, photon energy $\hbar\omega_{\text{pr}}=1.55\,\text{eV}$, pulse duration $\Delta\tau_{\text{p}}=120\,\text{fs}$, fluence $F_{\text{p}}=2.8\,\text{mJ}\,\text{cm}^{-2}$, repetition rate $f_{\text{rep}}=1\,\text{kHz}$) as a function of the time delay $\Delta t_{\text{p-pr}}$ (mechanical delay stage: Physik Instrumente M-414.DG), and for different combinations of incident and detected probe polarizations. Probe pulses are focused onto the sample in the $ac$-plane and at a $45^{\circ}$ angle to the surface normal using a plano-convex lens ($f_{\text{pump}}=75\,\text{mm}$). For the measurements shown in Figs. 1- 5, SHG signals are acquired in a cross-polarized configuration, with the incident (detected) probe polarization aligned parallel (perpendicular) to the $ac$-plane of the crystal.
@@
Parasitic SHG signals originating from optical elements in the probe path are suppressed using a long pass filter (Schott RG630, $2\,\text{mm}$). After reflection from the sample, second harmonic and fundamental beams are separated using a short pass filter (Schott BG39, $2\,\text{mm}$) and a bandpass filter (center wavelength $400\,\text{nm}$, bandwidth $25\,\text{nm}$). The SHG signal is collected by a plano-convex lens (focal length $f=75\,\text{mm}$), detected using a photomultiplier tube, and processed with electronic amplification and time-domain filtering via a boxcar integrator to enhance the signal-to-noise ratio. Each data point is averaged over 5000 pulses ($t_{\text{int}}=5$ s). The temporal overlap between pump and probe pulses (time zero) is determined by sum-frequency generation in a beta-barium-borate (BBO) crystal placed at the sample position. To ensure homogeneous excitation of the probed region, the pump spot size on the sample ($280\times 280\,\mu\text{m}$) is chosen significantly larger than the probe spot ($80\times 40\,\mu\text{m}$).
@@
For double-pulse trSHG measurements, pump-pulse pairs are generated using a Mach–Zehnder-type interferometer and focused colinearly onto the sample. The relative delay $\Delta t_{\text{p1-p2}}$ between pulses p1 and p2 is controlled by a motorized delay stage (Physik Instrumente M-111.DG) in one interferometer arm. The pulse energies and polarizations of p1 and p2 are adjusted independently using neutral density filters and half-wave plates placed in each arm. Temporal overlap ($\Delta t_{\text{p1-p2}}=0$) is determined by measuring the SHG signal from the sample as a function of $\Delta t_{\text{p1-p2}}$ for a fixed pump-probe delay ($\Delta t_{\text{p1-pr}}=100\,\text{ps}$) with both pump pulses polarized vertically relative to the $ac$-plane of the sample (see Fig. S1).
@@
To suppress coherent artifacts caused by interference of the two pump pulses near $\Delta t_{\text{p1-p2}}=0$, all actual measurements are performed with cross-polarized p1 and p2. Cross-correlation scans of the SHG signal as a function of the relative polarization angle $\alpha_{\text{p1,p2}}$ confirm the absence of measurable coherent artifacts at $\alpha_{\text{p1,p2}}=90^{\circ}$ (see Fig. S1). During double-pulse measurements, the SHG signal is recorded as a function of $\Delta t_{\text{p1-p2}}$ at each time point along the pump–probe delay, resulting in the two-dimensional data sets shown in Figs. 2c and 3a.
@@
Analysis of the pump-probe data. For the time-resolved polarizer scan in Fig. 1d, the SHG intensity $I_{\text{SHG}}(\Delta t_{\text{p-pr}}\,\phi)$ is normalized to the average pre-pump value $I_{\text{SHG}}(\Delta t_{\text{p-pr}}<0,\phi=90\,^{\circ})$. For the fluence- and delay-dependent traces in Fig. 1e, the SHG intensity $I_{\text{SHG}}(\Delta t_{\text{p-pr}})$ is normalized to the average pre-pump value $I_{\text{SHG}}(\Delta t_{\text{p-pr}}<0)$. For clarity, traces corresponding to different fluences are vertically offset by a constant. To extract the fluence dependent amplitudes, decay constants, sliding mode frequencies and long-lived signal suppression (see Fig. 1f and 1g), the model function
@@
is fitted to the delay-dependent traces in Fig. 1d. Here, $\Delta t_{\text{p-pr}}$ is the pump-probe delay, $\Delta I_{\text{PT}}$ corresponds to the change in the SHG signal due to the shift of the equilibrium position after photo-excitation (see also Ref. [21] for details), $\Delta I_{\text{SP}}$ is the modulation of the SHG intensity due to coherent sliding motion, and $\Delta I_{\text{fast}}^{\text{exc}}$ or $\Delta I_{\text{fast}}^{\text{decay}}$, respectively, model the fast initial suppression of the SHG signal. Parameters $\tau_{\text{PT}}$ and $\tau_{\text{sliding}}$ correspond to the characteristic time scale of the structural shift or the sliding mode coherence time, respectively. Furthermore, $\tau_{\text{fast}}^{\text{exc}}$ and $\tau_{\text{fast}}^{\text{decay}}$ are the time constants for the excitation and subsequent decay of the fast initial signal suppression observed in the $\chi_{zyy}$ component. $T_{\text{sliding}}$ is the sliding mode period.
Analysis of the pump-pump-probe data. To determine the sliding mode amplitude $A_{\text{s}}$, frequency $f_{\text{s}}$, phase $\phi_{\text{s}}$, and damping $\gamma_{\text{s}}=\tau_{\text{s}}^{-1}$, as well as the quasi-dc signal suppression $\Delta I_{\text{const}}$ in double-pulse experiments, we fit the model
@@
to the trSHG trace after the second excitation for each double-pulse delay $\Delta t_{\text{p-p}}$. The results are shown in Fig. 4a. The $\Delta t_{\text{p1-p2}}$-dependent sliding mode amplitude $A_{\text{s}}$ in Fig. 3f, on the other hand, is fitted by the model
@@
Here, $A_{\text{e}}^{\text{damp}}$ is the leverage of electronic excitation for the damping of the sliding mode amplitude and $\tau_{\text{e}}^{\text{damp}}$ the electronic relaxation time. To extract the Fourier amplitudes at frequencies $f_{\text{s}}$ and $2f_{\text{s}}$, we multiply the delay-dependent SHG trace after the second excitation pulse with a super-Gaussian window function
$F_{\text{filt},t}=\text{exp}\left(-\left(\frac{(\Delta t_{\text{p1-p2}}-t^{\prime})^{2}}{2\sigma_{t}}\right)^{3}\right),$ (6)
before applying a standard fast Fourier transform algorithm. The quotient $(\Delta\gamma_{\text{s}}/\Delta A_{\text{s}})(T)$ is determined by fitting the linear model
@@
Fitting of the polarization-dependent trSHG data. To fit the SHG polarimetry data shown in Fig. 1d, we consider the symmetry of the Td polytype. In the orthorhombic Td phase, WTe2 belongs to the $mm2$ point group. We consider only electric-dipole–allowed second-harmonic generation (SHG), described by the odd-parity, polar (i-type) second-order nonlinear susceptibility tensor $\chi^{(2)}$. According to Ref. [58], the non-vanishing components of $\chi^{(2)}$ for this symmetry are $\chi^{(2)}_{zzz}$, $\chi^{(2)}_{xxz}$, $\chi^{(2)}_{xzx}$, $\chi^{(2)}_{zxx}$, $\chi^{(2)}_{yyz}$, $\chi^{(2)}_{yzy}$, and $\chi^{(2)}_{zyy}$. The incident electric field is given by
@@
$I_{h}\sim|\hat{\bm{h}}\cdot\bm{P}|^{2}$ $=|\frac{1}{\sqrt{2}}(P_{x}+P_{z})|^{2}$ (12)
@@
Estimation of the transient lattice temperature increase. The transient temperature increase $\Delta T$ within the probed sample volume induced by the pump pulses in the pump–pump–probe experiments is estimated as
@@
Possible origin of the $2\omega_{\text{s}}$ component. The appearance of sidebands in the Fourier transform of the time-dependent SHG signal can be understood within the framework of the deformation potential formalism. Following the arguments of Fukuda et al. [37], who observed similar features in linear optical pump–probe experiments on Td-MoTe2, we consider an optical transition near resonance between valence and conduction band states. In the presence of a coherent phonon, the corresponding optical susceptibility is given by
$\chi_{0}(\omega)=\frac{|P_{\text{cv}}|^{2}}{\varepsilon_{0}}\frac{1}{\hbar(\omega-(\omega_{\text{cv}}-\Delta\omega_{\text{DP}})+\text{i}\Gamma}.$ (17)
Here, $\hbar\omega$ is the photon energy, $\hbar\omega_{\text{cv}}$ the energy difference between valence and conduction band states, $\hbar\Delta\omega_{\text{DP}}$ the phonon-induced energy shift, $\varepsilon_{0}$ the vacuum permittivity, $P_{\text{cv}}$ the optical transition matrix element, and $\Gamma$ accounts for damping. The deformation potential term can be expressed as $\hbar\Delta\omega_{\text{DP}}=h_{0}Q(t)$, where $h_{0}$ reflects the electron–phonon coupling strength and $Q(t)$ is the time-dependent phonon amplitude. Assuming $h_{0}Q(t)\ll\hbar(\omega-\omega_{\text{cv}})+\mathrm{i}\Gamma$, we expand $\chi(\omega)$ in powers of $Q(t)$:
$\chi(\omega,Q(t))$ $=\chi_{0}\left(1+\frac{h_{0}Q(t)}{\hbar(\omega-\omega_{\text{cv}})+\text{i}\Gamma}\right)^{-1}$ (18)
$\approx\chi_{0}\left(1+\left(\frac{h_{0}}{\hbar(\omega-\omega_{\text{cv}})+\text{i}\Gamma}\right)Q(t)+\left(\frac{h_{0}}{\hbar(\omega-\omega_{\text{cv}})+\text{i}\Gamma}\right)^{2}Q(t)^{2}\right).$ (19)
For $Q(t)\sim Q_{0}\cos(\omega_{\text{s}}t)$, $Q(t)^{2}\sim Q_{0}^{2}\cos^{2}(\omega_{\text{s}}t)=(Q_{0}^{2}/2)(1+\cos(2\omega_{\text{s}}t))$. With this, the second-order nonlinear polarization is given by
$P^{2\omega}_{i}$ $=\varepsilon_{0}\chi_{ijk}(\omega,Q(t))E_{j}E_{k}$ (20)
$=\varepsilon_{0}\chi_{ijk}^{0}(\omega)\left(1+\frac{Q_{0}^{2}\beta^{2}}{2}+Q_{0}\beta\cos(\omega_{\text{s}}t)+Q_{0}^{2}\beta^{2}\cos(2\omega_{\text{s}}t)\right)E_{j}E_{k},$ (21)
with $\beta=h_{0}/\hbar(\omega-\omega_{\text{cv}})+\text{i}\Gamma$, resulting in oscillations at frequency $2\omega_{\text{s}}$ of the SHG intensity in the time-domain.
Computational details. Density functional theory calculations are conducted within the plane-wave pseudopotential code Quantum ESPRESSO [59, 60], using fully-relativistic norm-conserving Pseudopotentials [61] and the Perdew-Burke-Ernzerhof generalized-gradient approximation (PBE) for the exchange-correlation functional [62]. The plane-wave kinetic-energy cutoff is set to 80 Ry, the Brillouin zone is sampled with a $10\times 6\times 4$ Monkhorst-Pack grid and the phonon dispersion is obtained from density functional perturbation theory on a $5\times 3\times 2$ q-point mesh [63]. The crystal lattice vectors are chosen to reproduce the experimental one from Ref. [64], and the crystal structure is relaxed within that unit cell. The electron and phonon eigenvalues, as well as the electron-phonon coupling matrix $g_{nm}^{\rm\nu}({\bf k},{\bf q})$, are interpolated on a dense $40\times 24\times 16$ grid via maximally-localized Wannier functions [65] within the EPW code [66], which uses Wannier90 as a library [67].
Ab initio description of coherent phonon excitation. The theoretical framework for the simulation of the coherent lattice dynamics is based on the solution of the coherent phonon equation of motion [38]:
@@
Here, $\partial_{t}=\partial/\partial_{t}$ and $Q_{\nu}$ is the adimensional displacement amplitude of the coherent phonon with frequency $\omega_{\nu}$. $\gamma_{\nu}$ is the coherent phonon damping rate mediated by phonon-phonon interactions [39]. The room-temperature coherent phonon damping rate is estimated from the third-order force constant, which is obtained from finite differences using the third-order.py utility of the shengBTE code [68]. The term on the right-hand side of Eq. 22 is the coherent phonon driving force due to electron-phonon interaction:

### v2

@@ Abstract
The sliding-mode frequency, on the other hand, exhibits a clear anti-correlation with the sliding amplitude (compare red and brown traces in Fig. 4a): as the amplitude increases, the frequency softens. Since both the lattice temperature and the level of electronic excitation are fixed in double-pulse experiments, a significant contribution from electronic softening, as reported in other semimetals [murray_effect_2005], can be excluded. We therefore attribute the observed amplitude-dependent frequency shift to intrinsic lattice anharmonicity, explored by large-amplitude coherent nuclear motion.
In pump–probe experiments, even such clear linear relations between coherent phonon amplitude, damping, and frequency are difficult to extract unambiguously, as variations in phonon amplitude are intertwined with electronic excitation. Double-pulse excitation overcomes this limitation by enabling independent control of the phonon amplitude within a targeted excited state.
@@
In summary, we have demonstrated that redistributing the energy of a single ultrafast excitation into two pulses overcomes the amplitude limitations of key vibrational modes, enhancing control over ferroic order and revealing nonequilibrium lattice dynamics. Our results show that the linear relationship between excited-carrier density and coherent vibrational amplitude assumed in DECP [zeiger_theory_1992] is an oversimplification, highlighting the potential of sequential excitation to surpass the resulting nonlinear limits and access large-amplitude dynamics. Such strategies are particularly advantageous in materials where decisive modes are exclusively Raman-active and coupled IR-active modes possess only weak dipole moments, rendering nonlinear phononic driving [forst_mode-selective_2015, nova_metastable_2019, disa_photo-induced_2023] ineffective.
Owing to the shared structural motifs of van der Waals materials and the universality of DECP, we expect the principles demonstrated here for WTe2 to be broadly applicable to other systems and their vibrational degrees of freedom, particularly other sliding ferroelectrics. In this context, large-amplitude vibrational spectroscopy enabled by sequential excitation provides a powerful means to investigate and tailor phonon-electron and phonon-phonon interactions far from equilibrium [cheng_coherent_2017, ravnik_time-domain_2021]. We envision that leveraging vibrational coherences near topologically trivial and paraelectric configurations opens new avenues for coherent control of both topology and ferroelectricity in functional materials.
@@ Methods
Sample preparation. WTe2 single crystals (commercially ordered from HQ Graphene) were exfoliated to remove the oxidized surface layers and immediately mounted in an optical-microscopy cryostat. The sample space was constantly purged with nitrogen gas to minimize oxidization of the pristine surface layers. SHG and oscillatory signals from the sample showed less than $2\,\%$ degradation over several weeks. All experiments were carried out at room temperature.
@@
Experimental setup. In single-pulse trSHG measurements, structural dynamics are initiated by focusing pump pulses (center wavelength $\lambda_{\text{p}}=1.5\,\mu\text{m}$, photon energy $\hbar\omega_{\text{p}}=0.83\,\text{eV}$, pulse duration $\Delta\tau_{\text{p}}=130\,\text{fs}$, absorbed fluence $F_{\text{p}}=0.4–6.5\,\text{mJ}\,\text{cm}^{-2}$, polarization along the $b$-axis of the WTe2 crystal, repetition rate $f_{\text{rep}}=1\,\text{kHz}$) onto the sample surface under normal incidence using a plano-convex lens ($f_{\text{pump}}=100\,\text{mm}$). The resulting dynamics are probed by monitoring the SHG signal produced by the probe pulses (center wavelength $\lambda_{\text{pr}}=0.8\,\mu\text{m}$, photon energy $\hbar\omega_{\text{pr}}=1.55\,\text{eV}$, pulse duration $\Delta\tau_{\text{p}}=120\,\text{fs}$, fluence $F_{\text{p}}=2.8\,\text{mJ}\,\text{cm}^{-2}$, repetition rate $f_{\text{rep}}=1\,\text{kHz}$) as a function of the time delay $\Delta t_{\text{p-pr}}$ set by a motorized delay stage, and for different combinations of incident and detected probe polarizations. Probe pulses are focused onto the sample in the $ac$-plane and at a $45^{\circ}$ angle to the surface normal using a plano-convex lens ($f_{\text{pump}}=75\,\text{mm}$). For the measurements shown in Figs. 1- 4, SHG signals are acquired in a cross-polarized configuration, with the incident (detected) probe polarization aligned parallel (perpendicular) to the $ac$-plane of the crystal.
@@
Parasitic SHG signals originating from optical elements in the probe path are suppressed using a long-pass coloured-glas filter (Schott RG630, $2\,\text{mm}$). After reflection from the sample, SHG and fundamental beams are separated using a short-pass filter (Schott BG39, $2\,\text{mm}$) and a band-pass filter (center wavelength $400\,\text{nm}$, bandwidth $25\,\text{nm}$). The SHG signal is collected by a plano-convex lens (focal length $f=75\,\text{mm}$), detected using a photomultiplier tube, and processed with electronic amplification and time-domain filtering via a boxcar integrator to enhance the signal-to-noise ratio. Each data point is averaged over 5000 pulses ($t_{\text{int}}=5$ s). The temporal overlap between pump and probe pulses (time zero) is determined by sum-frequency generation in a beta-barium-borate (BBO) crystal placed at the sample position. To ensure homogeneous excitation of the probed region, the pump spot size on the sample ($280\times 280\,\mu\text{m}^{2}$) is chosen significantly larger than the probe spot size ($80\times 40\,\mu\text{m}^{2}$).
@@
For double-pulse trSHG measurements, pump-pulse pairs are generated using a Mach–Zehnder-type interferometer and focused colinearly onto the sample. The relative delay $\Delta t_{\text{p1-p2}}$ between pulses p1 and p2 is controlled by a motorized delay stage in one interferometer arm. The pulse energies and polarizations of p1 and p2 are adjusted independently using neutral-density glass filters and half-wave plates placed in each arm. Temporal overlap ($\Delta t_{\text{p1-p2}}=0$) is determined by measuring the SHG signal from the sample as a function of $\Delta t_{\text{p1-p2}}$ for a fixed pump-probe delay ($\Delta t_{\text{p1-pr}}=100\,\text{ps}$) with both pump pulses polarized vertically relative to the $ac$-plane of the sample (see Fig. S1).
@@
To suppress coherent artifacts caused by interference of the two pump pulses near $\Delta t_{\text{p1-p2}}=0$, all actual measurements are performed with pump pulses p1 and p2 in a cross-polarized configuration. Cross-correlation scans of the SHG signal as a function of the relative polarization angle $\alpha_{\text{p1,p2}}$ confirm the absence of measurable coherent artifacts at $\alpha_{\text{p1,p2}}=90^{\circ}$ (see Fig. S1). During double-pulse measurements, the SHG signal is recorded as a function of $\Delta t_{\text{p1-p2}}$ at each time point along the pump–probe delay, resulting in the two-dimensional data sets shown in Figs. 2c and 3a.
@@
Analysis of the pump-probe data. For the time-resolved polarizer scan in Fig. 1c, the SHG intensity $I_{\text{SHG}}(\Delta t_{\text{p-pr}}\,\phi)$ is normalized to the average pre-pump value $I_{\text{SHG}}(\Delta t_{\text{p-pr}}<0,\phi=90\,^{\circ})$. For the fluence- and delay-dependent traces in Fig. 1e, the SHG intensity $I_{\text{SHG}}(\Delta t_{\text{p-pr}})$ is normalized to the average pre-pump value $I_{\text{SHG}}(\Delta t_{\text{p-pr}}<0)$. For clarity, traces corresponding to different fluences are vertically offset by a constant. To extract the fluence dependent amplitudes, decay constants, sliding-mode frequencies and long-lived signal suppression (see Fig. 1e and 1f), the model function
@@
is fitted to the delay-dependent traces in Fig. 1d. Here, $\Delta t_{\text{p-pr}}$ is the pump-probe delay, $\Delta I_{\text{PT}}$ corresponds to the change in the SHG signal due to the shift of the equilibrium position after photo-excitation (see also Ref. [sie_ultrafast_2019] for details), $\Delta I_{\text{SP}}$ is the modulation of the SHG intensity due to coherent sliding motion, and $\Delta I_{\text{fast}}^{\text{exc}}$ or $\Delta I_{\text{fast}}^{\text{decay}}$, respectively, model the fast initial suppression of the SHG signal. Parameters $\tau_{\text{PT}}$ and $\tau_{\text{sliding}}$ correspond to the characteristic time scale of the structural shift or the sliding mode coherence time, respectively. Furthermore, $\tau_{\text{fast}}^{\text{exc}}$ and $\tau_{\text{fast}}^{\text{decay}}$ are the time constants for the excitation and subsequent decay of the fast initial signal suppression observed in the $\chi_{zyy}$ component. $T_{\text{sliding}}$ is the sliding-mode period.
Analysis of the pump-pump-probe data. To determine the sliding-mode amplitude $A_{\text{s}}$, frequency $f_{\text{s}}$, phase $\phi_{\text{s}}$, and damping $\gamma_{\text{s}}=\tau_{\text{s}}^{-1}$, as well as the quasi-dc signal suppression $\Delta I_{\text{const}}$ in double-pulse experiments, we fit the model
@@
to the trSHG trace after the second excitation for each double-pulse delay $\Delta t_{\text{p-p}}$. The results are shown in Fig. 4a. The $\Delta t_{\text{p1-p2}}$-dependent sliding-mode amplitude $A_{\text{s}}$ in Fig. 3f, on the other hand, is fitted by the model
@@
Here, $A_{\text{e}}^{\text{damp}}$ is the leverage of electronic excitation for the damping of the sliding-mode amplitude and $\tau_{\text{e}}^{\text{damp}}$ the electronic relaxation time. Furthermore, the quotient $(\Delta\gamma_{\text{s}}/\Delta A_{\text{s}})(T)$ is determined by fitting the linear model
@@
Fitting of the polarization-dependent trSHG data. To fit the SHG polarimetry data shown in Fig. 1c and Fig. S2, we consider the symmetry of the Td polytype. In the orthorhombic Td phase, WTe2 belongs to the $mm2$ point group. We consider only electric-dipole–allowed SHG, described by the odd-parity, polar (i-type) second-order nonlinear susceptibility tensor $\chi^{(2)}$. According to Ref. [birss_macroscopic_1963], the non-vanishing components of $\chi^{(2)}$ for this symmetry are $\chi^{(2)}_{zzz}$, $\chi^{(2)}_{xxz}$, $\chi^{(2)}_{xzx}$, $\chi^{(2)}_{zxx}$, $\chi^{(2)}_{yyz}$, $\chi^{(2)}_{yzy}$, and $\chi^{(2)}_{zyy}$. The incident electric field is given by
@@
$I_{\text{h}}\sim|\hat{\bm{h}}\cdot\bm{P}|^{2}$ $=|\frac{1}{\sqrt{2}}(P_{x}+P_{z})|^{2}$ (11)
@@
While all accessible tensor elements oscillate coherently at the same frequency, a strong initial suppression is observed exclusively in the vertical lobe associated with the $\chi_{zyy}$ component (compare frames at $0.5\,\text{ps}$ and $0.5\,\text{ps}$ in Fig. 1c; for further details, see Fig. S2). We speculate that this anisotropic response results from an additional structural distortion along the $c$-axis, as suggested by density functional theory calculations under optical excitation [yang_laser-induced_2022].
Estimation of the transient lattice-temperature increase. The transient temperature increase $\Delta T$ within the probed sample volume induced by the pump pulses in the pump–pump–probe experiments is estimated as
@@
Computational details. Density functional theory calculations are conducted within the plane-wave pseudopotential code Quantum ESPRESSO [giannozzi_quantum_2009, giannozzi_advanced_2017], using fully-relativistic norm-conserving pseudopotentials [hamann_optimized_2013] and the Perdew-Burke-Ernzerhof generalized-gradient approximation (PBE) for the exchange-correlation functional [perdew_generalized_1996]. The plane-wave kinetic-energy cutoff is set to 80 Ry, the Brillouin zone is sampled with a $10\times 6\times 4$ Monkhorst-Pack grid, and the phonon dispersion is obtained from density functional perturbation theory on a $5\times 3\times 2$ q-point mesh [baroni_phonons_2001]. The crystal lattice vectors are chosen to reproduce the experimental ones from Ref. [mar_metal-metal_1992], and the crystal structure is relaxed within that unit cell. The electron and phonon eigenvalues, as well as the electron-phonon coupling matrix $g_{nm}^{\rm\nu}({\bf k},{\bf q})$, are interpolated on a dense $40\times 24\times 16$ grid via maximally-localized Wannier functions [marzari_maximally_2012] within the EPW code [lee_electronphonon_2023], which uses Wannier90 as a library [pizzi_wannier90_2020].
Ab initio description of coherent phonon excitation. The theoretical framework for the simulation of the coherent lattice dynamics is based on the solution of the coherent-phonon equation of motion [emeis_coherent_2025]:
@@
Here, $\partial_{t}=\partial/\partial_{t}$, and $Q_{\nu}$ is the adimensional displacement amplitude of the coherent phonon with frequency $\omega_{\nu}$. Also, $\gamma_{\nu}$ is the coherent-phonon damping rate mediated by phonon-phonon interactions [pan_origin_2025]. The room-temperature coherent phonon damping rate is estimated from the third-order force constant, which is obtained from finite differences using the third-order.py utility of the shengBTE code [li_shengbte_2014]. The term on the right-hand side of Eq. 16 is the coherent-phonon driving force due to electron-phonon interaction: